solve the following question with step by step.

∫∫ 2xy dxdy
1
Expert's answer
2020-10-13T15:23:24-0400

"\large\iint2xy\,dx\,dy" =


"\large\int(\int2xy\,dy)dx" =


"\large\int((2\times x\times\dfrac{y^{1+1}}{2})+c_1)dx" =


"\large\int (xy^2+c_1)dx" =


"\large\int xy^2+c_1\,\,dx" =


"\large\int xy^2\,dx+\int c_1\,dx" =


"\large(\int xy^2\,dx)+(c_1\int dx)" =


"\large(\dfrac{x^{1+1}}{2} \times y^2 ) +(c_1\times x) + c_2" =


"\large\dfrac{x^2y^2}{2} \, + c_1x+c_2"